### v1

@@ 3.2 Annihilate the unused part of the network
Proof 3.6 (Proof of Theorem 2.3).
For a given target function $f^{*}\in C([0,1])$, we first apply Theorem 2.2 to obtain a network $f^{\text{NN}}$ to approximate $f^{*}$, and then enhance the construction by pseudo-copy technique and constant-matching construction to eliminate the scaling $\gamma$ and shifting $\alpha$, leading to a network $f_{\text{pc}}^{\text{NN}}$ with fixed factors $\gamma=1,\alpha=0$.
a) Approximate the target function $f^{*}$ by a network $f^{\text{NN}}$ with learnable factors $\gamma$ and $\alpha$. By applying Theorem 2.2, we construct $f^{\text{NN}}$ to approximate $f^{*}$ through a piecewise constant function $g$ with error $\varepsilon$. Following the previous discussion, we have $\Delta h\leq\varepsilon/8$, $h=8L^{2}/(n-1)^{2}$ for some $L\in\mathbb{Z}^{+}$, and $\gamma=\Delta h/h=\Delta h(n-1)^{2}/8L^{2}$.
b) Remove the scaling $\gamma$ by the pseudo-copy technique. We first reassign the basis functions in $f^{\text{NN}}$ to construct the pseudo-copies of each approximator $f_{s_{j}}^{\text{NN}}$. To maintain the height the same as the original approximator, we choose $n=\sqrt{8L^{3}/\Delta h}+1$, such that $\gamma=L$. For each $K_{j}:=\{k_{j},k_{j}+L,k_{j}+2L,k_{j}+3L\}\subset K$, we denote the corresponding adjacent index set for pseudo-copy as $K_{j}^{\prime}=\{k_{j},k_{j}+1,\cdots,k_{j}+4L-1\}$ such that $K_{j}\subset K_{j}^{\prime}$. Here we partition $K_{j}^{\prime}$ into $L$ subsets as $K_{j}^{\prime}=\cup_{l=1}^{L}K_{j,l}$ by choosing after every $L$ indexes (illustrated at the top of Fig. 1(d)).
For each $l=1,\cdots,L$, the coefficients $\{p_{k},q_{k}\}_{k\in K_{j,l}}$ and {±$\{\pm b_{k}\}_{k\in K_{j,l}}$ are step-matching (according to Eq. (6)) to construct the pseudo-copy $f_{s_{j},p_{l}}^{\text{NN}}$ for $f_{s_{j}}^{\text{NN}}$ in Eq. (15), respectively. Since each $f_{s_{j},p_{l}}^{\text{NN}}$ has the same height $h$ with $f_{s_{j}}^{\text{NN}}$, we have the copy error as $|f_{s_{j},p_{l}}^{\text{NN}}(x)-f_{s_{j}}^{\text{NN}}(x)|<h$ for all $x\in[0,1]$. Therefore, Eq. (15) allows the summed $f_{s_{j},p_{l}}^{\text{NN}}$ to approximate $a_{j}f_{s_{j}}$ with the following error:
(18) $\begin{split}&\left|\sum_{l=1}^{L}\left[f_{s_{j},p_{l}}^{\text{NN}}(x)+a_{j}\frac{h}{2}\right]-a_{j}f_{s_{j}}(x)\right|\\ \leq\,&\sum_{l=1}^{L}\left|f_{s_{j},p_{l}}^{\text{NN}}(x)-f_{s_{j}}^{\text{NN}}(x)\right|+\left|L\left[f_{s_{j}}^{\text{NN}}(x)+a_{j}\frac{h}{2}\right]-a_{j}f_{s_{j}}(x)\right|\\ \leq\,&Lh+\Delta h=2\Delta h,\quad\forall\,x\in[0,1],\quad j=1,\cdots,J.\end{split}$
@@
c) Replace the shifting with constant-matching construction. Our constant-matching construction in Eq. (11) offers the necessary shifting $f_{c_{j}}^{\text{NN}}=a_{j}h/2$, enabling the combined $\sum_{l=1}^{L}(f_{s_{j},p_{l}}^{\text{NN}}+f_{c_{j}}^{\text{NN}})$ to approach $a_{j}f_{s_{j}}$. Denote $K_{j}^{c}$ for the index set used for the $L$-times construction of $f_{c_{j}}^{\text{NN}}$, and $K_{\text{use}}^{\prime}:=\cup_{j=1}^{J}(K_{j}^{p}\cup K_{j}^{c})\subset K$ as the index set for all involved basis functions. We define $\mathcal{S}_{\text{use}}^{\prime}$ as the picked subnetwork of $f_{\text{pc}}^{\text{NN}}$,
$\mathcal{S}_{\text{use}}^{\prime}(x):=\sum_{k\in K_{\text{use}}^{\prime}}\left[p_{k}\phi_{k}^{+}(x)+q_{k}\phi_{k}^{-}(x)\right]=\sum_{l=1}^{L}\sum_{j=1}^{J}\left[f_{s_{j},p_{l}}^{\text{NN}}(x)+f_{c_{j}}^{\text{NN}}(x)\right],\quad x\in[0,1].$
Eq. (18) implies the following error estimation,
(19) $\begin{split}E_{\text{use}}^{\prime}:=&\,\max_{x\in[0,1]}\left|\mathcal{S}_{\text{use}}^{\prime}(x)-g(x)\right|\\ =&\,\max_{x\in[0,1]}\left|\sum_{j=1}^{J}\sum_{l=1}^{L}\left[f_{s_{j},p_{l}}^{\text{NN}}(x)+f_{c_{j}}^{\text{NN}}(x)\right]-\sum_{j=1}^{J}a_{j}f_{s_{j}}(x)\right|\leq 2\Delta h.\end{split}$
@@
d) Eliminate the unused part of the network. Since the linear reorganization in Eq. (12) is invalid, we turn to apply the constant-matching construction in Eq. (11). to eliminate the unused part of the network. Concretely, for the remaining $n-8JL$ pairs of basis functions in $f_{\text{pc}}^{\text{NN}}$, we construct $f_{\pm c_{t}}^{\text{NN}}$, $t=1,\cdots,\lfloor n/8\rfloor-JL$, then pairing them up to offset with each other, i.e.,
$\mathcal{S_{\text{un}}^{\prime}}(x)=\sum_{k\in K\setminus K_{\text{use}}^{\prime}}\left[p_{k}\phi_{k}^{+}(x)+q_{k}\phi_{k}^{-}(x)\right]=\sum_{t=1}^{\lfloor n/8\rfloor-JL}\left[f_{c_{t}}^{\text{NN}}(x)+f_{-c_{t}}^{\text{NN}}(x)\right]+R(x).$
Here $R(x)$ is the residual part since $n$ is not always divisible by 8. However, our linear reorganization in Eq. (12) enables omit the basis functions with sufficiently small $b_{i}$.em without affecting the approximation result. Concretely, the $L^{\infty}$ error introduced by $R(x)$ can be controlled as $|R(x)|\leq\sum_{k=1}^{4}\bar{b}_{k}^{2}=:C_{R}$ for all $x\in[0,1]$, where $\{\bar{b}_{k}\}_{k=1}^{4}$ is the missed or external basis locations. Since the constant-matching construction has flexibility, $\{\bar{b}_{k}\}_{k=1}^{4}$ can be small enough to ensure $C_{R}\leq\Delta h$. Hence, the error $E_{\text{un}}^{\prime}$ introduced by the unused parameters satisfies
(20) $E_{\text{un}}^{\prime}=\max_{x\in[0,1]}\left|\mathcal{S_{\text{un}}^{\prime}}(x)\right|=C_{R}+C_{c}\leq 2\Delta h.$
@@
$\begin{split}\left|f_{\text{pc}}^{\text{NN}}(x)-g(x)\right|&=\left|\sum_{k\in K}\left[p_{k}\phi_{k}^{+}(x)+q_{k}\phi_{k}^{-}(x)\right]-g(x)\right|\\ &\leq\left|\mathcal{S}_{\text{use}}^{\prime}(x)-g(x)\right|+\left|\mathcal{S}_{\text{un}}^{\prime}(x)\right|\\ &\leq E_{\text{use}}^{\prime}+E_{\text{un}}^{\prime}\leq 4\Delta h,\quad\forall\,x\in[0,1],\end{split}$
where $\Delta h$ is chosen to satisfy $\Delta h\leq\varepsilon/8$. Hence we can finish the proof of Theorem 2.3 since $|f_{\text{pc}}^{\text{NN}}(x)-f^{*}(x)|\leq\varepsilon$ for all $x\in[0,1]$.
@@ 3.5 Estimate the approximation rate
Here we estimate the approximation rate of the $L^{2}$ error $e_{s}$ of approximating single step function $f_{s}$ in Eq. (3) by the approximator $f_{s}^{\text{NN}}$ in Eq. (7). In our four-pair construction, we assume $s=(b_{2}+b_{3})/2$ and introduce $k_{1}$ and $k_{2}$ to rewrite the symmetry relations in Eq. (4) as:
(21) $\begin{split}b_{1}=s-k_{2},\quad b_{3}=s+k_{1},\\ b_{2}=s-k_{1},\quad b_{4}=s+k_{2},\end{split}$
@@
(22) $\begin{split}e_{s}^{2}=&\int_{0}^{1}\left|\gamma\left[f_{s}^{\text{NN}}(x)+\frac{h}{2}\right]-f_{s}(x)\right|^{2}\mathop{}\!\mathrm{d}x\\ =&\,\gamma^{2}\left[\int_{b_{1}}^{s}\left|f_{s}^{\text{NN}}(x)+\frac{h}{2}\right|^{2}\mathop{}\!\mathrm{d}x+\int_{s}^{b_{4}}\left|f_{s}^{\text{NN}}(x)-\frac{h}{2}\right|^{2}\mathop{}\!\mathrm{d}x\right].\end{split}$
@@
(23) $\begin{split}e_{s}^{2}=\gamma^{2}\left[\frac{8}{3}(k_{1}-k_{2})^{2}\left(k_{1}^{3}+3k_{1}^{2}k_{2}+2k_{1}k_{2}^{2}+k_{2}^{3}\right)\right]\leq\gamma^{2}\frac{56}{3}d^{2}k_{2}^{3}.\end{split}$
@@
(24) $e_{s,p}=\left\|\sum_{l=1}^{L}\left(f_{s,p_{l}}^{\text{NN}}+\frac{h}{2}\right)-f_{s}\right\|_{L^{2}}\leq\sum_{l=1}^{L}\left\|\left(f_{s,p_{l}}^{\text{NN}}+\frac{h}{2M}\right)-\frac{1}{L}f_{s}\right\|_{L^{2}}=:\sum_{l=1}^{L}e_{s,p_{l}}.$
@@
(25) $\begin{split}\tilde{b}_{1}=(s+\Delta s_{l})-k_{2},\quad\tilde{b}_{3}=(s+\Delta s_{l})+k_{1},\\ \tilde{b}_{2}=(s+\Delta s_{l})-k_{1},\quad\tilde{b}_{4}=(s+\Delta s_{l})+k_{2}.\end{split}$
@@
(26) $\begin{split}e_{s,p_{l}}^{2}=&\int_{0}^{1}\left|\left[f_{s_{i}}^{\text{NN}}(x-\Delta s_{l})+\frac{h}{2}\right]-\frac{1}{L}f_{s}(x)\right|^{2}\mathop{}\!\mathrm{d}x\\ =&\int_{b_{1}+\Delta s_{l}}^{s}\left|f_{s_{i}}^{\text{NN}}(x-\Delta s_{l})+\frac{h}{2}\right|^{2}\mathop{}\!\mathrm{d}x+\int_{s}^{b_{4}+\Delta s_{l}}\left|f_{s_{i}}^{\text{NN}}(x-\Delta s_{l})-\frac{h}{2}\right|^{2}\mathop{}\!\mathrm{d}x,\end{split}$
@@
Therefore, we can first apply the UAP in the equidistant case to obtain a network $f_{\text{equi}}^{\text{NN}}$ in Eq. (2), which exhibits approximation power. Then, within a randomly initialized network $f_{\text{rand}}^{\text{NN}}$ of sufficient width, we find a subnetwork $f_{\text{sub}}^{\text{NN}}$ that can be regarded as randomly perturbed from $f_{\text{equi}}^{\text{NN}}$. If this perturbation is small enough, the subnetwork $f_{\text{sub}}^{\text{NN}}$ will also possess approximation power.
Proof 3.7 (Proof of Theorem 2.4).
@@
a) Approach $f^{*}$ with equidistantly initialized $f_{\text{equi}}^{\text{NN}}$. Theorem 2.2 indicate that for any small $\varepsilon$ and the target function $f^{*}$, there is an equidistantly initialized network $f_{\text{equi}}^{\text{NN}}$ in Eq. (2) with width $\hat{n}$ and $B_{\text{equi}}^{(\hat{n})}=(b_{i})_{i=1}^{\hat{n}}$, $W_{\text{equi}}^{(2\hat{n})}=(\pm b_{i})_{i=1}^{\hat{n}}$, such that
(27) $\left|f_{\text{equi}}^{\text{NN}}(x)-f^{*}(x)\right|<\varepsilon/4,\quad\forall\,x\in[0,1].$
b) Find a subnetwork $f_{\text{sub}}^{\text{NN}}$ in a randomly initialized $f_{\text{rand}}^{\text{NN}}$ to approximate $f_{\text{equi}}^{\text{NN}}$. For a network with sufficiently wide $n\gg\hat{n}$, parameters $B_{\text{rand}}^{(n)}\sim\mathcal{U}[0,1]^{n}$ and $W_{\text{rand}}^{(2n)}=(\!\pm p_{i})_{i=1}^{n}$, $p_{i}\sim\mathcal{U}[0,1]$, we can find a subnetwork $f_{\text{sub}}^{\text{NN}}$ with parameters
$B_{\text{sub}}^{(\hat{n})}=\left(b_{i}+r_{i}^{B}\right)_{i=1}^{\hat{n}},\quad W_{\text{sub}}^{(2\hat{n})}=\big{(}\pm\left(b_{i}+r_{i}^{W}\right)\big{)}_{i=1}^{2\hat{n}},$
which can be viewed as randomly perturbed from $B_{\text{equi}}^{(\hat{n})}$, $W_{\text{equi}}^{(2\hat{n})}$, while the independent and identically distributed (i.i.d.) $r_{i}^{B}\sim\mathcal{U}[-\Delta r,\Delta r]^{\hat{n}}$ and $r_{i}^{W}\sim\mathcal{U}[-\Delta r,\Delta r]^{2\hat{n}}$ measured the perturbation, and $\Delta r>0$ are the maximum allowable perturbation of $B_{\text{equi}}^{(\hat{n})}$ and $W_{\text{equi}}^{(2\hat{n})}$ (we impose further constraints on the parameters near the boundary of $[0,1]$ to prevent them from exceeding the range). Consequently, for sufficiently small $\Delta r<r_{0}$, the subnetwork $f_{\text{sub}}^{\text{NN}}$ will approach $f_{\text{equi}}^{\text{NN}}$ with the approximation error like
(28) $\left|f_{\text{sub}}^{\text{NN}}(x)-f_{\text{equi}}^{\text{NN}}(x)\right|<\varepsilon/4,\quad\forall\,x\in[0,1],\quad\text{ with probability }P_{\text{sub}},$

### v2

@@ 3.2 Annihilate the unused part of the network
Proof of Theorem 2.2.
For a given target function $f^{*}\in C([0,1])$, we first apply Theorem 2.1 to obtain a network $f^{\text{NN}}$ to approximate $f^{*}$, and then enhance the construction by pseudo-copy technique and constant-matching construction to eliminate the scaling $\gamma$ and shifting $\alpha$, leading to a network $f^{\text{NN}}$ in Eq. (2) with fixed factors $\gamma=1,\alpha=0$. To facilitate distinction, we will denote this network equipped with pseudo-copy technique as $f_{\text{pc}}^{\text{NN}}$ in the following text.
a) Approximate the target function $f^{*}$ by a network $f^{\text{NN}}$ with learnable factors $\gamma$ and $\alpha$. By applying Theorem 2.1, we construct $f^{\text{NN}}$ to approximate $f^{*}$ through a piecewise constant function $g$ with error $\varepsilon$. Following the previous discussion, we have $\Delta h\leq\varepsilon/8$, $\gamma=\Delta h/h$. During the construction, $L$ are chosen to be $L=\gamma$.
b) Remove the scaling $\gamma$ by the pseudo-copy technique. We first reassign the basis functions in $f^{\text{NN}}$ to construct the pseudo-copies of each approximator $f_{s_{j}}^{\text{NN}}$. For each $K_{j}=\{k_{j},k_{j}+L,k_{j}+2L,k_{j}+3L\}\subset K$, we denote the corresponding adjacent index set for pseudo-copy as $K_{j}^{\prime}:=\{k_{j},k_{j}+1,\cdots,k_{j}+4L-1\}$ such that $K_{j}\subset K_{j}^{\prime}$. To maintain the same height of each pseudo-copy, we partition $K_{j}^{\prime}$ into $L$ subsets as $K_{j}^{\prime}=\cup_{l=1}^{L}K_{j,l}$ by choosing after every $L$ indexes (illustrated at the top of Fig. 1(d)).
For each $l=1,\cdots,L$, we apply the step-matching construction on $\{p_{k},q_{k}\}_{k\in K_{j,l}}$ to construct the pseudo-copy $f_{s_{j},p_{l}}^{\text{NN}}$ for $f_{s_{j}}^{\text{NN}}$ in Eq. (15), respectively. Since each $f_{s_{j},p_{l}}^{\text{NN}}$ has the same height $h$ with $f_{s_{j}}^{\text{NN}}$, we have the copy error as $|f_{s_{j},p_{l}}^{\text{NN}}(x)-f_{s_{j}}^{\text{NN}}(x)|<h$ for all $x\in[0,1]$. Therefore, Eq. (15) allows the summed $f_{s_{j},p_{l}}^{\text{NN}}$ to approximate $a_{j}f_{s_{j}}$ with the following error:
$\begin{split}&\left|\sum_{l=1}^{L}\left[f_{s_{j},p_{l}}^{\text{NN}}(x)+a_{j}\frac{h}{2}\right]-a_{j}f_{s_{j}}(x)\right|\\ \leq\,&\sum_{l=1}^{L}\left|f_{s_{j},p_{l}}^{\text{NN}}(x)-f_{s_{j}}^{\text{NN}}(x)\right|+\left|L\left[f_{s_{j}}^{\text{NN}}(x)+a_{j}\frac{h}{2}\right]-a_{j}f_{s_{j}}(x)\right|\\ \leq\,&Lh+\Delta h=2\Delta h,\quad\forall\,x\in[0,1],\quad j=1,\cdots,J.\end{split}$ (18)
@@
c) Replace the shifting with constant-matching construction. After constructing the pseudo-copies $f_{s_{j},p_{l}}^{\text{NN}}$, our constant-matching construction in Eq. (11) can pair them up with the necessary shifting $f_{c_{j}}^{\text{NN}}=a_{j}h/2$, enabling the combined $\sum_{l=1}^{L}(f_{s_{j},p_{l}}^{\text{NN}}+f_{c_{j}}^{\text{NN}})$ to approach $a_{j}f_{s_{j}}$. Denote $K_{\text{use}}^{\prime}$ as the index set for all involved basis functions. We define $\mathcal{S}_{\text{use}}^{\prime}$ as the picked subnetwork of $f_{\text{pc}}^{\text{NN}}$,
$\mathcal{S}_{\text{use}}^{\prime}(x):=\sum_{k\in K_{\text{use}}^{\prime}}\left[p_{k}\phi_{k}^{+}(x)+q_{k}\phi_{k}^{-}(x)\right]=\sum_{l=1}^{L}\sum_{j=1}^{J}\left[f_{s_{j},p_{l}}^{\text{NN}}(x)+f_{c_{j}}^{\text{NN}}(x)\right].$
where $x\in[0,1]$. Eq. (18) implies the following error estimation,
$\begin{split}E_{\text{use}}^{\prime}:=&\,\max_{x\in[0,1]}\left|\mathcal{S}_{\text{use}}^{\prime}(x)-g(x)\right|\\ =&\,\max_{x\in[0,1]}\left|\sum_{j=1}^{J}\sum_{l=1}^{L}\left[f_{s_{j},p_{l}}^{\text{NN}}(x)+f_{c_{j}}^{\text{NN}}(x)\right]-\sum_{j=1}^{J}a_{j}f_{s_{j}}(x)\right|\leq 2\Delta h.\end{split}$ (19)
@@
d) Eliminate the unused part of the network. Since the linear reorganization in Eq. (12) is invalid, we turn to apply the constant-matching construction in Eq. (11) to eliminate the unused part of the network. Concretely, for the remaining $n-8JL$ pairs of basis functions in $f_{\text{pc}}^{\text{NN}}$, we construct $f_{\pm c_{t}}^{\text{NN}}$, $t=1,\cdots,\lfloor n/8\rfloor-JL$, then pairing them up to offset with each other, i.e.,
$\mathcal{S_{\text{un}}^{\prime}}(x)=\sum_{k\in K\setminus K_{\text{use}}^{\prime}}\left[p_{k}\phi_{k}^{+}(x)+q_{k}\phi_{k}^{-}(x)\right]=\sum_{t=1}^{\lfloor n/8\rfloor-JL}\left[f_{c_{t}}^{\text{NN}}(x)+f_{-c_{t}}^{\text{NN}}(x)\right]+R(x),$
Here $R(x)$ is the residual part since $n$ is not always divisible by 8. However, our linear reorganization in Eq. (12) enables omit the basis functions with sufficiently small $b_{i}$. Concretely, the $L^{\infty}$ error introduced by $R(x)$ can be controlled as $|R(x)|\leq\sum_{k=1}^{4}\bar{b}_{k}^{2}=:C_{R}$ for all $x\in[0,1]$, where $\{\bar{b}_{k}\}_{k=1}^{4}$ is the missed or external basis locations. Since the constant-matching construction has flexibility, $\{\bar{b}_{k}\}_{k=1}^{4}$ can be small enough to ensure $C_{R}\leq\Delta h$. Hence, the error $E_{\text{un}}^{\prime}$ introduced by the unused parameters satisfies
$E_{\text{un}}^{\prime}=\max_{x\in[0,1]}\left|\mathcal{S_{\text{un}}^{\prime}}(x)\right|=C_{R}+C_{c}\leq 2\Delta h.$ (20)
@@
$\begin{split}\left|f_{\text{pc}}^{\text{NN}}(x)-g(x)\right|&=\left|\sum_{k\in K}\left[p_{k}\phi_{k}^{+}(x)+q_{k}\phi_{k}^{-}(x)\right]-g(x)\right|\\ &\leq\left|\mathcal{S}_{\text{use}}^{\prime}(x)-g(x)\right|+\left|\mathcal{S}_{\text{un}}^{\prime}(x)\right|\\ &\leq E_{\text{use}}^{\prime}+E_{\text{un}}^{\prime}\leq 4\Delta h,\quad\forall\,x\in[0,1].\end{split}$
where $\Delta h$ is chosen to satisfy $\Delta h\leq\varepsilon/8$. Hence we can finish the proof of Theorem 2.2 since $|f_{\text{pc}}^{\text{NN}}(x)-f^{*}(x)|\leq\varepsilon$ for all $x\in[0,1]$. ∎
@@ 3.5 Estimate the approximation rate
Here we estimate the approximation rate of the $L^{2}$-error $e_{s}$ of approximating single step function $f_{s}$ in Eq. (3) by the approximator $f_{s}^{\text{NN}}$ in Eq. (7). In our four-pair construction, we assume $s=(b_{2}+b_{3})/2$ and introduce $k_{1}$ and $k_{2}$ to rewrite the symmetry relations in Eq. (5) as:
$\begin{split}b_{1}=s-k_{2},\quad b_{3}=s+k_{1},\\ b_{2}=s-k_{1},\quad b_{4}=s+k_{2},\end{split}$ (21)
@@
$\begin{split}e_{s}^{2}=&\int_{0}^{1}\left|\gamma\left[f_{s}^{\text{NN}}(x)+\frac{h}{2}\right]-\frac{h}{\Delta h}f_{s}(x)\right|^{2}\mathop{}\!\mathrm{d}x\\ =&\,\gamma^{2}\left[\int_{b_{1}}^{s}\left|f_{s}^{\text{NN}}(x)+\frac{h}{2}\right|^{2}\mathop{}\!\mathrm{d}x+\int_{s}^{b_{4}}\left|f_{s}^{\text{NN}}(x)-\frac{h}{2}\right|^{2}\mathop{}\!\mathrm{d}x\right].\end{split}$ (22)
@@
$\begin{split}e_{s}^{2}=\gamma^{2}\left[\frac{8}{3}(k_{1}-k_{2})^{2}\left(k_{1}^{3}+3k_{1}^{2}k_{2}+2k_{1}k_{2}^{2}+k_{2}^{3}\right)\right]\leq\gamma^{2}\frac{56}{3}d^{2}k_{2}^{3}.\end{split}$ (23)
@@
$e_{s,p}=\left\|\sum_{l=1}^{L}\left(f_{s,p_{l}}^{\text{NN}}+\frac{h}{2}\right)-f_{s}\right\|_{L^{2}}\leq\sum_{l=1}^{L}\left\|\left(f_{s,p_{l}}^{\text{NN}}+\frac{h}{2M}\right)-\frac{1}{L}f_{s}\right\|_{L^{2}}=:\sum_{l=1}^{L}e_{s,p_{l}}.$ (24)
@@
$\begin{split}\tilde{b}_{1}=(s+\Delta s_{l})-k_{2},\quad\tilde{b}_{3}=(s+\Delta s_{l})+k_{1},\\ \tilde{b}_{2}=(s+\Delta s_{l})-k_{1},\quad\tilde{b}_{4}=(s+\Delta s_{l})+k_{2}.\end{split}$ (25)
@@
$\begin{split}e_{s,p_{l}}^{2}=&\int_{0}^{1}\left|\left[f_{s_{i}}^{\text{NN}}(x-\Delta s_{l})+\frac{h}{2}\right]-\frac{1}{L}f_{s}(x)\right|^{2}\mathop{}\!\mathrm{d}x\\ =&\int_{b_{1}+\Delta s_{l}}^{s}\left|f_{s_{i}}^{\text{NN}}(x-\Delta s_{l})+\frac{h}{2}\right|^{2}\mathop{}\!\mathrm{d}x+\int_{s}^{b_{4}+\Delta s_{l}}\left|f_{s_{i}}^{\text{NN}}(x-\Delta s_{l})-\frac{h}{2}\right|^{2}\mathop{}\!\mathrm{d}x,\end{split}$ (26)
@@
Therefore, we can first apply the UAP in the equidistant case to obtain a network $f^{\text{NN}}$ in Eq. (2) (denoted as $f_{\text{equi}}^{\text{NN}}$ in the following text for distinction), which exhibits approximation power. Then, within a randomly initialized network $f^{\text{NN}}$ in Eq. (2) (denoted as $f_{\text{rand}}^{\text{NN}}$) of sufficient width, we find a subnetwork $f_{\text{sub}}^{\text{NN}}$ that can be regarded as randomly perturbed from $f_{\text{equi}}^{\text{NN}}$. If this perturbation is small enough, the subnetwork $f_{\text{sub}}^{\text{NN}}$ will also possess approximation power.
Proof of Theorem 2.3.
@@
a) Approach $f^{*}$ with equidistantly initialized $f_{\text{equi}}^{\text{NN}}$. Theorem 2.1 indicate that for any small $\varepsilon$ and the target function $f^{*}$, there is an equidistantly initialized network $f_{\text{equi}}^{\text{NN}}$ in Eq. (2) with width $\tilde{n}$ and $B_{\text{equi}}^{(\tilde{n})}=(b_{i})_{i=1}^{\tilde{n}}$, $W_{\text{equi}}^{(2\tilde{n})}=(\pm b_{i})_{i=1}^{\tilde{n}}$, such that
$\left|f_{\text{equi}}^{\text{NN}}(x)-f^{*}(x)\right|<\varepsilon/4,\quad\forall\,x\in[0,1].$ (27)
b) Find a subnetwork $f_{\text{sub}}^{\text{NN}}$ in a randomly initialized $f_{\text{rand}}^{\text{NN}}$ to approximate $f_{\text{equi}}^{\text{NN}}$. For a network with sufficiently wide $n\gg\tilde{n}$, parameters $B_{\text{rand}}^{(n)}\sim\mathcal{U}[0,1]^{n}$ and $W_{\text{rand}}^{(2n)}=(\!\pm p_{i})_{i=1}^{n}$, $p_{i}\sim\mathcal{U}[0,1]$, we can find a subnetwork $f_{\text{sub}}^{\text{NN}}$ with parameters
$B_{\text{sub}}^{(\tilde{n})}=\left(b_{i}+r_{i}^{B}\right)_{i=1}^{\tilde{n}},\quad W_{\text{sub}}^{(2\tilde{n})}=\big{(}\pm\left(b_{i}+r_{i}^{W}\right)\big{)}_{i=1}^{2\tilde{n}},$
which can be viewed as randomly perturbed from $B_{\text{equi}}^{(\tilde{n})}$, $W_{\text{equi}}^{(2\tilde{n})}$, while the independent and identically distributed (i.i.d.) $r_{i}^{B}\sim\mathcal{U}[-\Delta r,\Delta r]^{\tilde{n}}$ and $r_{i}^{W}\sim\mathcal{U}[-\Delta r,\Delta r]^{2\tilde{n}}$ measured the perturbation, and $\Delta r>0$ are the maximum allowable perturbation of $B_{\text{equi}}^{(\tilde{n})}$ and $W_{\text{equi}}^{(2\tilde{n})}$ we impose further constraints on the parameters near the boundary of $[0,1]$ to prevent them from exceeding the range). Consequently, for sufficiently small $\Delta r<r_{0}$, the subnetwork $f_{\text{sub}}^{\text{NN}}$ will approach $f_{\text{equi}}^{\text{NN}}$ with the approximation error like
$\left|f_{\text{sub}}^{\text{NN}}(x)-f_{\text{equi}}^{\text{NN}}(x)\right|<\varepsilon/4,\quad\forall\,x\in[0,1],\quad\text{ with probability }P_{\text{sub}}.$ (28)